The wheel of a bicycle of radius 90 cm is moving with the speed of 50 km per hour. Calculate its angular speed in seconds. (Show your calculations)
Answer by Alan3354(69443)

The wheel of a bicycle of radius 90 cm is moving with the speed of 50 km per hour. Calculate its angular speed in seconds.
----
50 km/hr = 5000000 cm/hr = 12500/9 cm/sec
---
The wheel moves 180*pi cm per revolution.
(12500/9)/(180*pi) = 625/(81*pi) rev/sec
625/(81*pi) rev/sec **** angular speed in rev/sec
= 625*2*pi/(81*pi) = 1250/81 radians/sec **** angular speed in rads/sec